# Root Defence Strategies: Ensuring Safety of LLM at the Decoding Level

### v1

@@ Abstract
Large language models (LLMs) have demonstrated immense utility across various industries. However, as LLMs advance, the risk of harmful outputs increases due to incorrect or malicious instruction prompts. While current methods effectively address jailbreak risks, they share common limitations: 1) Judging harmful responses from the prefill-level lacks utilization of the model’s decoding outputs, leading to relatively lower effectiveness and robustness. 2) Rejecting potentially harmful responses based on a single evaluation can significantly impair the model’s helpfulness. This paper examines the LLMs’ capability to recognize harmful outputs, revealing and quantifying their proficiency in assessing the danger of previous tokens. Motivated by pilot experiment results, we design a robust defense mechanism at the decoding level. Our novel decoder-oriented, step-by-step defense architecture corrects harmful queries directly rather than rejecting them outright. We introduce speculative decoding to enhance usability and facilitate deployment to boost secure decoding speed. Extensive experiments demonstrate that our approach improves model security without compromising reasoning speed. Notably, our method leverages the model’s ability to discern hazardous information, maintaining its helpfulness compared to existing methods.
@@ 1 Introduction
In recent years, significant progress has been made in developing large language models (LLMs). Meanwhile, the safety of LLMs has attracted significant attention from the research community and industry (Weidinger et al., 2021; Achiam et al., 2023; Wu et al., 2023b). One of the primary safety concerns is jailbreaking, where malicious actors or errant inputs prompt LLMs to produce harmful or inappropriate content, effectively bypassing ethical guidelines. Many attempts have been made to address these risks. For example, Meta has implemented several strategies in both pre-training and fine-tuning phases to improve the safety of their Llama-series models (Touvron et al., 2023; Dubey et al., 2024). Despite these efforts, some studies have reported that focusing too narrowly on safety may diminish the models’ general capability (Bai et al., 2022; Huang et al., 2024). Therefore, enhancing LLMs’ safety without compromising their utility has become a critical area of research.
@@
Recent defense strategies against jailbreaks can be roughly categorized into two groups (as shown in Figure 1). The first group is prefill-level defense (Wu et al., 2023a; Phute et al., 2023; Zheng et al., 2024). It enhances the models’ protective capabilities by integrating additional security measures into the initial prompts (prefills) or refining their representation. However, this approach primarily depends on user inputs to detect harmful responses, making it susceptible to rapidly advancing jailbreaking techniques. Moreover, this reliance can lead to inaccuracies in interpreting user intentions, thereby reducing the overall utility of the LLMs. Another group of methods is response-level defenses (Phute et al., 2023; Xu et al., 2024). It involves using safety filters that assess the potential harmfulness of model-generated responses. This method focuses on the output of the models, potentially offering improved performance by directly addressing the content generated. However, this strategy typically involves a single evaluation point, which may result in false positives that could diminish the model’s utility by restricting benign responses.
@@
Figure 1: Examples of two kinds of imperfect defenses and RDS. (a) Prefill-level defenses fail to refuse the harmful query with $N$ harmful tokens. (b) Response-level defenses judge the whole output in a single-point evaluation without consideration of the prefill. (c) RDS conducts step-by-step assessments for each sampled token to enhance the security of LLMs at the decoder level.
@@
In practice, jailbreak instructions can bypass the prefill-level defenses and achieve their purposes in the model’s response Wei et al. (2024). Therefore, assessing jailbreak behavior in LLMs should focus on decoding dimensions, including the context of both the prefill and the model’s response. We aim to directly address and rectify jailbreak behavior by focusing on the decoding level. (Zheng et al., 2024) has demonstrated models’ ability to distinguish between harmful and benign prefill. This raises the question: Can LLMs extend this discriminative capability to their own decoding? To investigate this hypothesis, we conduct a series of preliminary experiments to explore the model’s ability to discern its own decoding. Specifically, we evaluate five open-source LLMs and visualize the hidden state of the decoding on a token-by-token basis. We observe that LLMs cannot distinguish harmful tokens from harmless tokens in one step, but it can achieve identification through multi-step judgment at the decoding, especially for harmful prefill.
@@
Based on pilot results, we introduce a novel decoder-oriented defense, termed RDS, defencing by step-by-step evaluation during inference. Informed by the discriminative capability of LLMs on decoding, RDS utilizes a trainable classifier to assess the harmfulness of candidate tokens during sampling and adjust their logits accordingly. Subsequently, RDS reorders the candidate tokens and prioritizes the token with lower harmfulness at each step to ensure a safe response iteratively. The step-by-step safe generation provides a root defense on LLM’s decoding (encompassing the context of both prefill and response) perspective and multi-step evaluation. Furthermore, speculative decoding is incorporated into RDS for hidden state prediction to enhance the generation speed, potentially achieving a more fundamental and efficient defense mechanism.
@@
We evaluate RDS on five LLMs and a series of harmful and harmless query benchmarks. Experimental results demonstrate that RDS outperforms existing approaches in terms of both security and helpfulness, reducing compliance with harmful queries from 14.4% to 2.4% on Xstest ((Röttger et al., 2023)) (without safety prompt) and increasing token generation speed by $\mathbf{2.12\times\sim 3.09\times}$ compared to other baselines. We hope this method offers a new perspective to security defense, i.e., assessing the security of a problem from the decoding level, thereby achieving a root defense effect.
2 Preliminary: Decoding-level Defence
@@ 2.1 LLMs’ Discriminative Capability of Decoding
The prefill stage for LLMs typically includes a user query, often accompanied by prefixed or suffixed elements such as system prompts. Previous studies have demonstrated that LLMs can discriminate between different types of prefill and use this ability to enhance safety mechanisms. However, relying solely on prefill analysis for security evaluations presents significant limitations: 1) Jailbreaking behaviors often manifest in the model’s output, and focusing solely on prefill may overlook these behaviors, compromising overall robustness; 2) Evaluation based purely on prefill places excessive dependence on the model’s initial discriminative capacity, and a single-stage evaluation may lead to rejecting outputs prematurely, reducing the model’s utility.
@@
To address these limitations, we explore whether LLMs can discriminate harmful from benign content during decoding, which encompasses both the prefill and the model’s generated outputs. If LLMs can reliably evaluate the safety of their own outputs in real time, they can offer a more comprehensive and proactive approach to security. Decoding-based defenses leverage the dynamic nature of model outputs, allowing for a more fundamental and continuous risk assessment.
2.2 Experiments
@@
In this section, we evaluate five open-source LLMs and utilize Principal Component Analysis (PCA) to visualize their hidden states during the decoding process. To facilitate classifier training, we curate Custom from the DRO (Zheng et al., 2024) as the training dataset of the classifier, consisting of 100 harmful and 100 benign queries, all beginning with the phrase “How to”. The evaluated LLMs are accessible on HuggingFace: Llama-2-chat-7B (Touvron et al., 2023), Llama-3-8b-Instruct (AI@Meta, 2024), Qwen2-7B-Instruct (Yang et al., 2024), Vicuna-7B-v1.3, and Vicuna-13B-v1.3 (Chiang et al., 2023). Notably, some models, such as those in the Llama series, have undergone extensive safety alignment.
Following previous studies (Zheng et al., 2024), we use the hidden states $\mathbf{h}$ of the prefill at the top layer of the model for classifier training. Details of the classifier’s training objective is provided in Appendix C. The classifier is applied to the decoding sampling strategy. Refusal responses often start with special tokens, such as “I’m sorry” or “As an AI”, which distinguish them from compliant responses. Given the distinct nature of refusal and compliant responses, particularly at the start of the output, we focus on the first three tokens and the last token of the output (i.e., $i\in{1,2,3,N}$, where $N$ is the length of the output), visualizing the hidden states of each token from the top layer.
@@
The responses to harmful queries are assessed using Llama-guard (Bhatt et al., 2023), which is a safety classification model based on LLaMA-2 (Touvron et al., 2023). Benign responses are evaluated through string matching. If refusal strings are identified in the response, it is categorized as a refusal response; otherwise, it is not. A compliant answer is assigned an evaluation score $s$ of 1, otherwise 0. Lower scores are preferable for harmful queries, while higher scores are ideal for benign queries. In the preliminary experiment, we sample one response for each query. The initial defense of these five LLMs is presented in Appendix D.
@@ 2.3 Visualization Analysis
We apply PCA to visualize the hidden state and compute the first four principal components of the hidden states. Figure 2 respectively show the visual results from the first to the third token and the last token of the refusal responses for harmful queries (represented by “refusal+i”) and compliance for benign queries (represented by “compliance+i”). See the visual results from the 4-th token to the 7-th in Appendix E. In Figure 2, harmful and benign queries can be clearly distinguished at the hidden state layer. The boundary, depicted by a black dashed line, separates these two queries using training with only a few parameters. LLMs naturally can discern the harmfulness of inputs.
Can LLMs extend this discriminative capability to their own decoding? In Figure 2, from 1-th to 3-th token, almost all the tokens to benign queries maintain at the benign side for all LLMs. However, for harmful queries, although refusal tokens refer to safe responses, the first few tokens are not classified correctly. Instead, we observe that the refusal tokens to harmful queries tend to converge towards the benign side with a smaller offset than harmful tokens. This shift is most pronounced in llama-2-chat-7B and Qwen2-7B-Instruct, whose responses to harmful queries are all refusal. Even on models with poor defense, this shift still exists. We interpret these observations as the discriminative capacity of LLMs to their own decoding.
Can we consider LLMs’ output harmful based on a single judgment? Figure 2 illustrates that the classifier cannot accurately determine whether the output is harmful based solely on the model’s overall decoding (i.e., the complete output). Even current advanced methods cannot guarantee 100% filtering. Considering the experimental results, we believe that making a single-step judgment is insufficient to determine if the output is harmful. In conjunction with Figure 2, although the model cannot make an accurate judgment in one attempt, it can achieve better discrimination through a step-by-step evaluation of the decoding. Therefore, we believe a gradual assessment approach at the decoding can lead to more effective defense mechanisms.
@@
(4) i=last
Figure 2: Performance of the classifier at the decoding from the 1-th to 3-th token and last token of the output. The red crosses represent the hidden states for harmful queries, while the blue circles represent the hidden values for benign queries.
@@ 3 Methodology
Motivated by validating the capability to recognize responses, we propose RDS to ensure the safety of LLMs at the decoder level. The architecture of RDS is illustrated in Figure 3. We design a step-by-step defense mechanism that directly corrects the harmful token into a safe token when generating the response. Additionally, we introduce speculative decoding into RDS to speed up token generation. Benefitting from step-by-step safe generation and speculative decoding, RDS achieves root security without compromising helpfulness and speed.
Figure 3: RDS comprises two key modules: 1) Step-by-step token generation: The root classifier is designed based on the discriminative capacity of queries. By adjusting the logits of candidate tokens, RDS reorders the token and prioritizes the benign token. 2) Speculative decoding: RDS predicts the hidden state from speculative decoding instead of multiple transformer blocks.
@@ 3.1 Problem Formulation
Let $x_{i}$ as the model’s decoding at step $t_{i}$, $c_{i}=f(x_{i})$ represents the score of the sampled token $x_{i}$ calculated by the classifier $f(\cdot)$, where the training object of the classifier is illustrated in Appendix C. RDS aims to minimize $c_{i}$ at each step. We formulate this process as follows:
$\underset{x_{i}}{\min}\sum_{i=1}^{N}c_{i}\quad;\quad x_{i}=\text{LLM}(x_{i-1};\mathbb{C}_{i})$ (1)
where $N$ is the length of outputs, and at each step $t_{i}$, the LLM obtains prior decoding $x_{i-1}$ and the harmful results $\mathbb{C}_{i}$ of candidate tokens to generate next token $x_{i}$. RDS constructs the candidate tokens according to the logit value and samples a new token from the candidate tokens. By ensuring the security of each step, RDS promises a safe response.
@@ 3.2 Step-by-step safe generation
where $\mathbf{l}_{i-1}=\text{LM\_Head}(\mathbf{h}_{i-1})$ represents the logits of the decoding at step $t_{i-1}$, $\mathbf{h}_{i-1}$ represents the hidden state of the decoding at step $t_{i-1}$, $\mathbb{I}_{i}$ and $\mathbb{V}_{i}$ represent the set of top-$k$ candidate tokens and the logits values of these candidate tokens, respectively.
@@
3.3 Speculative decoding
RDS leverages the discriminative ability of decoding for defense by computing the harmful score of candidate tokens based on their hidden states. It concatenates decoding at step $t-1$ with candidate tokens to obtain the hidden state at step $t$ resembling speculative decoding processes (such as EAGLE (Li et al., 2024)) that predict hidden states from decoding and tokens. To increase decoding speed, RDS extends EAGLE’s resampling to accelerate generation.

### v2

@@ Abstract
Large language models (LLMs) have demonstrated immense utility across various industries. However, as LLMs advance, the risk of harmful outputs increases due to incorrect or malicious prompts. While current methods effectively address jailbreak risks, they share common limitations: 1) Judging harmful outputs from the prefill-level lacks utilization of the model’s decoding outputs, leading to relatively lower effectiveness and robustness. 2) Rejecting potentially harmful outputs based on a single evaluation can significantly impair the model’s helpfulness. To address the above issues, we examine LLMs’ capability to recognize harmful outputs, revealing and quantifying their proficiency in assessing the danger of previous tokens. Motivated by pilot experiment results, we design a robust defense mechanism at the decoding level. Our novel decoder-oriented, step-by-step defense architecture corrects the outputs of harmful queries directly rather than rejecting them outright. We introduce speculative decoding to enhance usability and facilitate deployment to boost safe decoding speed. Extensive experiments demonstrate that our approach improves model security without compromising reasoning speed. Notably, our method leverages the model’s ability to discern hazardous information, maintaining its helpfulness compared to existing methods.
Root Defence Strategies: Ensuring Safety of LLM at the Decoding Level
Xinyi Zeng1thanks: Equal contribution, Yuying Shang [^1] footnotemark: 1, Jiawei Chen3, Jingyuan Zhang4,Yu Tian2thanks: Corresponding author 1Aerospace Information Research Institute, Chinese Academy of Sciences, Beijing, China 2Dept. of Comp. Sci. and Tech., Institute for AI, Tsinghua University, Beijing, China 3Shanghai Key Laboratory of Multi. Info. Processing, East China Normal University 4Kuaishou-Inc tianyu1810613@gmail.com
@@ 1 Introduction
Large language models (LLMs) have advanced significantly in recent years, prompting growing attention from academia and industry to their safety implications (Weidinger et al., 2021; Achiam et al., 2023; Wu et al., 2023b). One of the primary safety concerns is jailbreaking, where malicious actors or errant inputs prompt LLMs to produce harmful or inappropriate content, effectively bypassing ethical guidelines. Many attempts have been made to address these risks. For instance, Meta has implemented several strategies in both pre-training and fine-tuning phases to improve the safety of their Llama-series models (Touvron et al., 2023; Dubey et al., 2024). Despite these efforts, some studies have reported that focusing too narrowly on safety may diminish the models’ general capability (Bai et al., 2022; Huang et al., 2024). Therefore, enhancing LLMs’ safety without compromising their utility has become a critical area of research.
@@
Recent defense strategies against jailbreaks can be roughly categorized into two groups (as shown in Figure 1). The first group is prefill-level defense (Wu et al., 2023a; Phute et al., 2023; Zheng et al., 2024). It enhances the models’ protective capabilities by integrating additional security measures into the initial prompts (prefills) or refining their representation. However, this approach primarily depends on user inputs to detect harmful outputs, making it susceptible to rapidly advancing jailbreaking techniques. Moreover, this reliance can lead to inaccuracies in interpreting user intentions, thereby reducing the overall utility of the LLMs. Another group of methods is output-level defenses (Phute et al., 2023; Xu et al., 2024). It involves using safety filters that assess the potential harmfulness of model-generated outputs. This method focuses on the output of LLMs, potentially offering improved performance by directly addressing the content generated. However, this strategy typically involves a single evaluation point, which may result in false positives that could diminish the model’s utility by restricting benign outputs.
@@
Figure 1: Examples of recenqt imperfect defenses and RDS. a) Prefill-level defenses fail to refuse the harmful query with $N$ harmful tokens. b) Output-level defenses judge the whole output in a single-point evaluation without consideration of the prefill. c) RDS conducts step-by-step assessments for each sampled token to enhance the security of LLMs at the decoder level.
@@
In practice, jailbreak instructions can bypass the prefill-level defenses and achieve their purposes in the model’s output Wei et al. (2024). Therefore, assessing jailbreak behavior in LLMs should focus on decoding dimensions, including the context of both the prefill and the model’s output. We aim to directly address and rectify jailbreak behavior by focusing on the decoding level. (Zheng et al., 2024) has demonstrated models’ ability to distinguish between harmful and benign prefill. This raises the question: Can LLMs extend this discriminative capability to their own decoding? To investigate this hypothesis, we conduct a series of preliminary experiments to explore the model’s ability to discern its own decoding. Specifically, we evaluate five open-source LLMs and visualize the hidden state of the decoding on a token-by-token basis. We observe that LLMs cannot distinguish harmful tokens from benign tokens in one step, but it can achieve identification through multi-step judgment at the decoding, especially for harmful prefill.
@@
Based on pilot experiment results, we introduce a novel decoder-oriented defense, termed RDS, defending by step-by-step evaluation. Informed by the discriminative capability of LLMs on decoding, RDS utilizes a trainable classifier to assess the harmfulness of candidate tokens during sampling and prioritizes the token with lower harmfulness at each step to ensure a safe output iteratively. The step-by-step safe generation provides a root defense on LLM’s decoding (encompassing the context of both prefill and output) perspective and multi-step evaluation. Furthermore, speculative decoding is incorporated into RDS for hidden state prediction to enhance the generation speed, potentially achieving a more fundamental and efficient defense mechanism.
@@
We evaluate RDS on five LLMs and a series of harmful and benign query benchmarks. Experimental results demonstrate that RDS outperforms existing approaches in terms of both security and helpfulness, reducing compliance with harmful queries from 2.0% to 37% and increasing token generation speed by $\mathbf{2.12\times\sim 3.09\times}$. We hope this method offers a new perspective to security defense, i.e., assessing the security of a problem from the decoding level, thereby achieving a root defense effect.
2 Preliminary: Decoding-level Defense
@@ 2.1 LLMs’ Discriminative Capability of Decoding
The prefill stage for LLMs typically includes a user query, often accompanied by prefixed or suffixed elements such as system prompts. Previous study (Zheng et al., 2024) has demonstrated that LLMs can discriminate between different types of prefill and use this ability to enhance safety mechanisms. However, solely relying on prefill analysis for security evaluations presents significant limitations: 1) Jailbreaking behaviors often manifest in the model’s output, and focusing solely on prefill may overlook these behaviors, compromising overall robustness; 2) Evaluation based purely on prefill places excessive dependence on the model’s initial discriminative capacity, and a single-stage evaluation may lead to rejecting outputs prematurely, reducing the model’s utility.
@@
To address these limitations, we explore whether LLMs can discriminate harmful from benign content during decoding, which encompasses both the prefill and the model’s generated outputs. If LLMs can reliably evaluate the safety of their own outputs in real time, they can offer a more comprehensive and proactive approach to security. Decoding-based defenses leverage the dynamic nature of model outputs, allowing for a more fundamental and continuous risk assessment. We use the hidden states of the harmful and benign queries at the top layer of the model for classifier training. Details of the classifier’s training objective is provided as follows.
$\mathbf{u}=\frac{1}{n}{\textstyle\sum_{q=1}^{n}}\mathbf{h}^{q},$ (1)
$\mathbf{m_{i}}={\mathbf{V}}^{T}(\mathbf{h_{i}}-\mathbf{u}),$ (2)
$\hat{y_{i}}={\mathbf{W}}^{T}\mathbf{m_{i}}+\mathbf{b},$ (3)
$\mathcal{L}(y_{i},\hat{y}_{i})=-\frac{1}{n}{\textstyle\sum_{q=1}^{n}}(y_{i}\log{\hat{y}_{i}}+(1-y_{i})\log{(1-\hat{y}_{i}})),$ (4)
where $\mathbf{u}\in\mathbb{R}^{d}$ is the mean value of all hidden states of queries, $\mathbf{V}\in\mathbb{R}^{d\times m}$ represents the $m$ principal components, $\mathbf{W}\in\mathbb{R}^{1\times d}$ and $\mathbf{b}\in\mathbb{R}^{1}$ are the trainable parameters. $\hat{y}_{i}$ and $y_{i}$ represent the predicted score and the label of query, respectively.
2.2 Preliminary settings
@@
We utilize Principal Component Analysis (PCA) to visualize the hidden states during the decoding process. To facilitate classifier training, we curate the training dataset Custom from DRO Zheng et al. (2024) to fit the classifier, consisting of 100 harmful and 100 benign queries. The evaluated LLMs are: Llama-2-chat-7B (Touvron et al., 2023), Llama-3-8b-Instruct (AI@Meta, 2024), Qwen2-7B-Instruct (Yang et al., 2024), Vicuna-7B-v1.3, and Vicuna-13B-v1.3 (Chiang et al., 2023). Notably, some models, such as Llama-2-chat-7B, have been aligned in safety.
@@
We visualize the hidden state from the top layer of each generated token to verify the classifier ability at decoding. The outputs of harmful queries are assessed using Llama-guard (Bhatt et al., 2023), which is a safety classification model based on LLaMA-2 (Touvron et al., 2023). If refusal strings are identified in the output, it is categorized as a refusal output; otherwise, it is not. A compliant answer is assigned an evaluation score $s$ of 1, otherwise 0. The compliant outputs to harmful queries are treated as harmful outputs. Others including the refusal outputs to harmful queries and benign queries, and compliant outputs to benign queries are treated as benign outputs. In the preliminary experiment, we sample one output for each query. The initial defense of these five LLMs is presented in Appendix C.
@@ 2.3 Visualization Analysis
We apply PCA to visualize the hidden state and select the first four principal components of the hidden states. Refusal outputs often start with special tokens, such as “I’m sorry” or “As an AI”. As refusal outputs are distinguished from compliant outputs at the start, we samples the first few tokens to verify the classifier performance on output. Besides, we additionally sample the last token of the output. Figure 2 respectively show the visual results of the first eight tokens of the outputs. The boundary (the black dashed line) separates harmful queries (red cross) and benign queries (blue circles), which illustrates that LLMs can naturally discern the harmfulness of the inputs.
@@
(4) $i$=4
Figure 2: Performance of the classifier at the decoding from the $i$-th token of the output. Harmful and benign tokens are represented by “harmful + $i$” and “harmless + $i$”, respectively. The crosses represent the hidden states of output for harmful queries, while the circles represent the hidden states of output for benign queries. See the visual results from the 5-th token to the 8-th in Appendix D.
Can LLMs extend this discriminative capability to their own decoding? In Figure 2, from 1-th to 4-th token, almost all the tokens to benign queries maintain at the benign side. Although refusal tokens to harmful queries refer to benign outputs, some of them maintain at the harmful side. While compliant tokens maintain at the benign side. The classifier performs poorly in hard classification. On the contrary, we observe that benign tokens of harmful queries are closer to the harmful side compared to harmful tokens. That is to say, for harmful queries, benign tokens receive higher scores from the classifier than harmful tokens, which means a distribution differentiation rather than hard classification. We interpret the distribution differentiation between harmful and benign tokens as the LLMs’ discriminative capacity of LLMs of decoding.
Can LLMs recognize benign decoding based on a single judgment? The current step confirms the safety of the immediate decoding without guaranteeing the safety of subsequent decoding. Making a single-step judgment is insufficient to ensure the safety of whole output. Due to the random sampling strategy, we observe that there is a phenomenon of rejecting first and then answering in the outputs. As described in (Zhou et al., 2024), deepening the consistency of security measures beyond merely aligning the first few tokens can significantly improve the security of LLMs. Therefore, we believe a step-by-step assessment approach at the decoding can ensure the robustness of defense.
@@ 3 Methodology
Motivated by validating the capability to recognize outputs, we propose RDS to ensure the safety of LLMs at the decoder level. The architecture of RDS is illustrated in Figure 3. We design a step-by-step defense mechanism that directly corrects the harmful token into a safe token when generating the output. Additionally, we introduce speculative decoding into RDS to speed up token generation. Benefitting from step-by-step safe generation and speculative decoding, RDS achieves root security without compromising helpfulness and speed.
Figure 3: RDS comprises two key modules: 1) Step-by-step safe generation: The root classifier is designed based on the discriminative capacity of queries. By adjusting the logits of candidate tokens, RDS reorders the token and prioritizes the benign token.2) Hidden State Prediction: Based on the hidden state at step $t-1$ and candidate token embedding, RDS calculate the hidden state of candidate token from speculative head instead of multiple transformer blocks. The hidden state at step $t$ refers to the hidden state of the final selected token. $h_{0}$ represents the hidden state of the last token of queries.
@@ 3.1 Problem Formulation
Let $x_{i}$ as the LLM’s decoding at step $t_{i}$, the step-by-step inference process of LLMs can be formulates as: $\quad x_{i}=[x_{i-1};\max(\mathbb{V}_{i})]$, where $x_{0}$ represents the query, $\mathbb{V}_{i}$ represents the logits over vocabulary. RDS aims to ensure the safety of token sampling at each step, which can be formulates as:
$\quad x_{i}=[x_{i-1};\max(\mathbb{C}_{i})];\mathbb{C}_{i}=f(\mathbb{I}_{i},x_{i-1})$ (5)
where $N$ is the length of outputs, [;][;][ ; ] represents concatenate operation, $\mathbb{C}_{i}$ represents the score of candidate tokens calculated by the classifier $f(\cdot)$. By ensuring the security from step 1 to $N$, RDS promises a safe output.
@@ 3.2 Step-by-step safe generation
where $\mathbf{l}_{i-1}=\text{LM\_Head}(\mathbf{h}_{i-1})$ represents logits at step $t_{i-1}$, $\mathbf{h}_{i-1}$ represents the hidden state of the decoding at step $t_{i-1}$, $\mathbb{I}_{i}$ and $\mathbb{V}_{i}$ represent the set of top-$k$ candidate tokens and the logits values of these candidate tokens, respectively.
@@
3.3 Hidden State Prediction
RDS leverages the discriminative ability of decoding for defense by computing the harmful score of candidate tokens based on their hidden states. It concatenates decoding at step $t-1$ with candidate tokens to obtain the hidden state at step $t$ resembling EAGLE (Li et al., 2024) that predict hidden states from decoding and tokens. RDS extends EAGLE_Head in resampling process to generate the hidden state of the candidate tokens.